rth n S 26 lia c IGv A randonh selacted pregant bomcn A vandom ee prg rcand om
Answer #1

Total sample size = 30,300,000 + 500,000 + 3000 + 100 = 30,803,100

a)

Probability of delivering twins = 500,000 / 30,803,100 = 0.01623

b)

Probability of delivering quadruplets = 100 / 30,803,100 = 3.25 x 10-6 = 0.00000325

c)

Probability of delivering more than a single child = (500,000 + 3,000 + 100) / 30,803,100 = 503,100 / 30,803,100

= 0.01633

Cheers!

Upvote if this was helpful.